[figure: 9.7 cm]

MAGNETIC AND TRANSPORT PROPERTIES OF THIN FILMS

OF Gd_{2}Fe_{12}Cr_{2}B PREPARED BY FLASH-EVAPORATION

I. Gościańska^{a}, T Toliński^{b}, A. Kowalczyk^{b}, J. Baszyński^{b}, A. Szlaferek^{b}

aDepartment of Physics, A. Mickiewicz University, Poznań, Poland bInstitute of Molecular Physics, PAS, Smoluchowskiego 17, 60-179 Poznań, Poland

Gd_{2}Fe_{12}Cr_{2}B alloys are prepared in a shape of thin films by a flesh-evaporation technique. The samples are amorphous if prepared on a cooled glass substrate and crystallize after annealing. Samples obtained on a non-cooled substrate are micro- or nanocrystalline. Magnetic properties are studied by vibrating sample magnetometer and the temperature dependence of resistance is measured in the range 4.2 K÷800 K. The resistance of the amorphous films increases with decreasing the temperature from R.T. down to 4.2 K, while the annealed samples exhibit metallic behavior. The crystallization process was observed by resistance measurements in temperature varied from RT to 823 K at a rate of 10 K/min.

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : poster

Corresponding author : I. Gościańska

Address for correspondence : Department of Physics, A. Mickiewicz University, Poznań, Poland

Email address : igosc@amu.edu.pl